CHIRPED-PULSED FTMW SPECTRUM OF 4-FLUOROBENZYL ALCOHOL. STRUCTURE AND TOR-SIONAL MOTIONS IN THE GROUND ELECTRONIC STATE PHASE^{a}

\underline{RYAN G. BIRD} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, Pa 15213; JUSTIN L. NEILL and BROOKS H. PATE, Department of Chemistry, University of Virginia, Charlottesville, Va 22904.

Previous studies of 4-fluorobenzyl alcohol (4FBA) in the gas phase have revealed the existence of tunneling splittings in its high resolution electronic spectrum. These could be attributed either to the ground S_{0} state, or the excited S_{1} state, or both. Motivated by this finding, we report here a study of the CP-FTMW spectrum of 4FBA. Its pure rotational spectrum was collected using a series of 250 MHz chirped pulses, over the frequency range of 6.5-18 GHz. Analysis of this data shows that 4FBA has a gauche structure and yields the V_{2} barrier for CH_{2}OH torsion, from which the predicted tunneling splittings in the ground state have been extracted.

aWork supported by NSF (CHE-0618740 and -0911117).